OpenStudy (anonymous):

Explain in your own word how factoring is used to solve quadratic equations. Demonstrate the process with an example.

OpenStudy (anonymous):

x^2+x-2=0 (x-1)(x+2) we set each factor to equal zero and solve for x: x-1=0, which means x=1 and x+2=0 which means x=-2